LBC & LUCIFER Preparations for Partner Science:
UT 2010 Apr 07

Observer: OKuhn, MPedani (Tucson), DThompson (LBTO)
Telescope Operator: David Gonzalez Huerta
Instrument Support: TShih
Telescope Support: JLittle
SW support: Dan Cox (on call)

Summary

  • Primary instrument used will be LBC
  • Closed for winds between 6:00-8:30
  • Switch to LUCIFER at 09:30, but ran into faults
    • M2 went haywire, required power cycle on crate
    • M3 not in simulation mode
  • Corrected pointing: IA=-384.4
  • Took 2 pairs of images for orientation check
  • Checked collimation model - looks good.

Preparations

Details

UT 00:15 LUCIFER UI started as observer
- The telescope-IIF connection was down. TCS was running, so I restarted the telescope service (00:24) in the Control SW panel as engineer. Dark 2s image taken, looks normal. Confirmed image was saved to /data/luci/20100407 and the archive.

After reviewing logs from last several nights, it is not clear the field aberrations are correctly set for LUCIFER. So, we will switch to LUCIFER as soon as some remaining tests from LBC are completed, and gather a set of data with LUCIFER.

09:29 DGH is switching to LUCIFER.

Set up LUCIFER: BrGam+clear, N3.75, No mask in use. Files: luci.20100407.NNNN.fits

0002: N1.8 camera
0003: N3.75 camera...check if positioned correctly. It looks OK, the gap is even on both sides, and a little smaller along the bottom vs. at the top. The rounded corners are visible in all 4 corners, but cut off slightly on the bottom edge.
0004: I moved back to the N1.8, initialized, and then back to the N3.75. The camera went back to the same position, so I will leave it here and continue.

10:04 Initial presets had problems (no AGw selected, then M3 not homed [but why is it active?], giving an "open loop" error message). Initial pointing OK (~24 arcsec off in CA) but after clearing all active optics the initial collimation was quite good.

10:20 Called Dan Cox, 10:25 called Paul Grenz, no answer. 10:32 called Norm. 11:10 have to move to horizon and physically reset something on M2. DGH & JLittle cycled power on the M2 crate, this seems to have fixed M2. JL confirms that M3 is disconnected. It should be in simulation mode but apparently is not...Norm to correct this in LBT.conf.

11:36 Re-opening dome.

11:40 Preset to WT10_249 (EL 83), WT10_053 (EL 9), IA=-384.4. Back to WT10_249, minor correction to CA.

11:58 Preset to BS9170 to collimate. #0013 30.0s in BrGam filter, #0014 offset 15,15 for skysub (use for orientation) Tout=-2.1C.

12:13 Preset to BS9178 #0017, #0018 2x15s (use for orientation). Guider image elongated. EL~50, wfs rms ~165 at frame 0019.

Check collimation on pointing model stars, track mode presets:
0021 EL=60
0022 EL=49
0023 EL=40
0024 EL=27
0025 EL=15
0026 EL=60
0027 EL=75
0028 EL=80
0029 EL=86
Collimation model looks good over whole range, pointing too.

12:40 Closing dome.

Observing Plan (LUCIFER)

  • Closed dome SW tests, if requested

  • Twilight flats in NB, BB filters

  • Correct initial collimation/pointing:
    • Use Brackett Gamma filter, can do in twilight
    • Correct IE/CA at high elevation
    • Correct IA at low elevation (AZ encoder problems)
    • Check/correct IE/CA again at high elevation

  • Open dome SW tests, if requested

  • Run dice5 pattern to check that things still work
    • Check repeated positioning on LUCIFER with and without flex comp on

  • Gather NB illumination correction data

  • Re-test updatePointingReference functionality.

  • Repeat measurement of focal plane scale, orientation

  • Check existing collimation model
    • Take new data if necessary
    • Stable conditions!

  • Check existing pointing model
    • Implement new collimation model if necessary
    • Stable conditions!
    • Take 12 points, with pointing log on...
    • Take full set (56 points) if the first 12 points are off.

  • Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
    • Check to make sure offsets still happen in a reasonable manner

  • Collect through-focus sequence for Andrew - reasonably dense star field.
    • also check the AGw-LUCIFER focus in cold weather for temp term?

  • Collect astrometric field data to check AGw calibration at colder temps

  • Improve map of AGw patrol field limits at high field angle
    • Requires set up work on a stone field
    • See this plot for extra patrol field area (in blue): AGw_patrolfield_v2.pdf

  • Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

-- DavidThompson - 07 Apr 2010
Topic revision: r3 - 07 Apr 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback